OpenStudy (anonymous):

I need some help on a few problems for my quiz tomorrow, so stick around and explain if you can, please! Solve the equation by finding square roots. 9x^2+12x+4=5

OpenStudy (anonymous):

subtract the 5 so you get -1 9x^2+12x-1=0 *can't factor so quadratic equation* \[-12\sqrt{12^2 -4(9)(-1)}/2(9)\] solve the rest

OpenStudy (anonymous):

We've never done it like that before in class, so I'm honestly lost.

OpenStudy (anonymous):

what do you do in class

OpenStudy (anonymous):

We normally would do something like subtract the 4 and get 1 on the other side, then figure out (#x+#)^2 = 5 then solve from there, but I'm not exactly sure how to get the two numbers I substituted with #.

OpenStudy (anonymous):

I'm really not sure, this is all completely new to me.

OpenStudy (anonymous):

(#x+#)^@ = 1, I mean, but tbh I don't really know what I'm saying at all now

OpenStudy (anonymous):

you can try this since i have no idea what you wrote

OpenStudy (anonymous):

I have no idea what you wrote either, so hopefully there will be someone else that learned the useless way I did.........